OpenStudy (anonymous):

y=lmx->inf (xe^x)/(x^2+2x-1) use l hopitals rule

zepdrix (zepdrix):

Before using L'Hopital's Rule, verify that the limit is of an indeterminate form. This limit appears to be of the indeterminate form: \[\frac{ \infty }{ \infty }\] So we can apply L'Hopital's Rule. Do you need help applying the rule? :)

OpenStudy (anonymous):

yeah i do pls

OpenStudy (anonymous):

beacuse after applying the rule, i syill got inf/inf

zepdrix (zepdrix):

 Still getting Inf/Inf? Good that is still an indeterminate form, which means you can apply L'Hop again! :)